# Confidence Calibration and Rationalization for LLMs via Multi-Agent Deliberation

### v1

@@ 1 Introduction
Although confidence estimation and calibration have been extensively studied in the broader machine learning literature (Gal & Ghahramani, 2016; Guo et al., 2017), previous work in the context of NLP mostly required extensive fine-tuning (Kong et al., 2020) or temperature-based scaling (Guo et al., 2017; Jiang et al., 2021), which can be expensive for LLMs. Recent studies on confidence estimation for black-box LLMs adopted either consistency-based approaches with repeated sampling (Wang et al., 2023b) or verbalization-based elicitation through direct prompting (Lin et al., 2022; Tian et al., 2023), or combined (Xiong et al., 2023). However, results are mixed on whether the elicited confidences are better-calibrated than sample-based estimates or the model’s original token probabilities, and no attempt has been made to further improve the interpretability and reliability of such measurement by incorporating the rationalization and collective refinement of individual confidence, leveraging the expressive, self-critical, tool-use and collaborative capabilities of generative LLM agents (Park et al., 2023; Madaan et al., 2023; Schick et al., 2023).
@@ 2 Related Work
Confidence Estimation and Calibration for LLMs: With the rapid advancement of LLMs, there has been an increasing focus on estimating and calibrating their prediction confidence, primarily for classification settings (Jiang et al., 2021; Si et al., 2022; 2023a; Portillo Wightman et al., 2023). For the more challenging yet pertinent tasks of free-form generation with varying answer lengths, Kuhn et al. (2023) introduce an unsupervised entropy-based metric that captures the uncertainty over meanings rather than sequences, and Liu et al. (2023b) present a lightweight training method that learns a bias term added to the output logits for better-calibrated confidence estimates. However, these methods require access to internal model structures or the output logits, which are not available for proprietary LLMs. Recent research endeavors have focused on either estimating confidence based on the probability distribution of the most consistent answer from multiple samples (Wang et al., 2023b; Chen et al., 2023d; Lin et al., 2023) or directly eliciting verbalized confidence (Lin et al., 2022; Tian et al., 2023), and found that prompting or fine-tuning with verbalized expression of uncertainty can lead to increased accuracy or better calibration (Mielke et al., 2022; Zhou et al., 2023). Xiong et al. (2023) present a comprehensive analysis of logit-based, verbalized, and consistency-based methods and suggest that a hybrid approach incorporating verbalized confidence into a consistency-based ensemble is more effective for confidence calibration.
@@
LLM Agent Ensemble: Augmented with external memory system and tool-use abilities, LLM-powered language agents (Park et al., 2023) display great potential to serve as self-consistent human proxies that can plan by self-reflection and refinement (Shinn et al., 2023; Madaan et al., 2023; Sun et al., 2023). Harnessing the diversity and collaboration strengths of multiple LLM agents, the ensemble and interaction can effectively enhance their capabilities in complex reasoning (Xiong et al., 2023; Chen et al., 2023b; Wang et al., 2023a), instruction following (Jiang et al., 2023b), and value alignment (Liu et al., 2023a). Du et al. (2023) introduce a multi-agent debate framework where multiple LLM instances propose and debate their reasoning processes to reach a consensus, which improves the factuality and performance on arithmetic and strategic reasoning tasks. Similarly, ReConcile (Chen et al., 2023a) uses different backbone LLMs for debating and a confidence-weighted voting mechanism to obtain the final answer. Our approach differs in that we incorporate self-consistency estimates, and naturally calibrate the verbalized confidence through group deliberation with peer feedback and intermediate rationales for confidence adjustment, rather than manual rescaling in a post-hoc manner. We also incorporate both open and API-based models for the flexibility of supporting both logit-based and black-box confidence estimation. Liu et al. (2023c) propose an inference-time agent selection framework based on aggregated peer ratings over multiple rounds of interactions for a given query. In contrast, our agent selection process requires only the first-round individual confidence estimates at the task level.
@@ 3.1 Agent Ensemble and Stance Generation
To strive for a balance between task accuracy and calibration performance, it is necessary to maintain a certain level of diversity among the agents’ initial answers and reasoning paths while allocating the slots wisely so that the most suitable agents for the task can ideally become the majority, which serves as the basis for the ensemble. Inspired by Si et al. (2023b), we initialize four types of ”expert agents”. Each agent has a different prompting strategy or tool-use expertise: Chain-of-Thought (Wei et al., 2022) for multi-hop reasoning tasks, Program-of-Thoughts (Chen et al., 2023c) for code and arithmetic reasoning tasks, Search-Augmented Self-Ask (Press et al., 2023) for factoid reasoning tasks, and GenRead prompting (Yu et al., 2023) for knowledge-intensive reasoning tasks. This initialization is flexible in that any new skill or prompting strategy can be easily added with modularity. As multiple skills might be relevant for an input dataset, we determine the importance ranking of each skill and accordingly allocate expert agents, based on a simple uncertainty-aware calibration score, detailed in Appendix A.1. We use Mistral-7B (Jiang et al., 2023a), GPT-3.5-turbo (OpenAI, 2022), and Cohere-Commend (Cohere, 2023) as backbones for the expert agents. For an incoming test question (e.g., ”Which element was discovered in 1898 and named after Greek ’new’?”), each expert agent performs self-deliberation by executing the corresponding prompting strategy and voting independently for an answer with a confidence estimate. The numerical confidence estimate can be based on the output sequence perplexity (PP) for open models  [^2] For open models with access to next-token probabilities, the raw confidence for the output sequence $W=(w_{1},...,w_{N})$ can be estimated as $\frac{1}{PP(W)}=P(w_{1},...,w_{N})^{1/N}$., or verbally elicited for black-box models, normalized to $[0,1]$. Note that this estimate is often unreliable and needs further calibration. We then obtain a set of unique and diverse answers (”stances”) by merging semantically equivalent pairs into clusters, using a GPT-3.5 judge following Tian et al. (2023). This constitutes the output of the first stage: semantically unique stances each with a corresponding frequency and aggregated mean confidence.
@@ A.4 Prompt templates and example output
We follow the prompt templates from the official implementations of CoT (Wei et al., 2022), PoT (Chen et al., 2023c), and Knowledge prompting (Yu et al., 2023). For cost considerations, we adopt the zero-shot setting for each strategy (e.g. ”Let’s think/write Python programs step-by-step” without providing a demonstration) and resort to search verification only when the model is unsure about a generated premise. Other prompt templates, including the prompts for argument generation, feedback, and final confidence can be found in Table 2.

### v3

@@ 1 Introduction
Although confidence estimation and calibration have been extensively studied in the broader machine learning literature (Gal & Ghahramani, 2016; Guo et al., 2017), previous work in the context of NLP mostly required extensive fine-tuning (Kong et al., 2020) or temperature-based scaling (Guo et al., 2017; Jiang et al., 2021), which can be expensive for LLMs. Recent studies on confidence estimation for black-box LLMs adopted either consistency-based approaches with repeated sampling (Wang et al., 2023b) or verbalization-based elicitation through direct prompting (Lin et al., 2022; Tian et al., 2023), or combined (Xiong et al., 2023). However, results are mixed on whether the elicited confidences are better-calibrated than sample-based estimates or the model’s original token probabilities, and no attempt has been made to further improve the interpretability and reliability of such measurement by incorporating the rationalization and collective refinement of individual confidence, leveraging the expressive, self-critical, tool-use and collaborative capabilities of generative LLM agents (Park et al., 2023; Madaan et al., 2024; Schick et al., 2024).
@@ 2 Related Work
Confidence Estimation and Calibration for LLMs: With the rapid advancement of LLMs, there has been an increasing focus on estimating and calibrating their prediction confidence, primarily for classification settings (Jiang et al., 2021; Si et al., 2022; 2023a; Portillo Wightman et al., 2023). For the more challenging yet pertinent tasks of free-form generation with varying answer lengths, Kuhn et al. (2023) introduce an unsupervised entropy-based metric that captures the uncertainty over meanings rather than sequences, and Liu et al. (2024) present a lightweight training method that learns a bias term added to the output logits for better-calibrated confidence estimates. However, these methods require access to internal model structures or the output logits, which are not available for proprietary LLMs. Recent research endeavors have focused on either estimating confidence based on the probability distribution of the most consistent answer from multiple samples (Wang et al., 2023b; Chen et al., 2023c; Lin et al., 2023) or directly eliciting verbalized confidence (Lin et al., 2022; Tian et al., 2023), and found that prompting or fine-tuning with verbalized expression of uncertainty can lead to increased accuracy or better calibration (Mielke et al., 2022; Zhou et al., 2023). Xiong et al. (2023) present a comprehensive analysis of logit-based, verbalized, and consistency-based methods and suggest that a hybrid approach incorporating verbalized confidence into a consistency-based ensemble is more effective for confidence calibration.
@@
LLM Agent Ensemble: Augmented with external memory system and tool-use abilities, LLM-powered language agents (Park et al., 2023) display great potential to serve as self-consistent human proxies that can plan by self-reflection and refinement (Shinn et al., 2024; Madaan et al., 2024; Sun et al., 2024). Harnessing the diversity and collaboration strengths of multiple LLM agents, the ensemble and interaction can effectively enhance their capabilities in complex reasoning (Xiong et al., 2023; Chen et al., 2023b; Wang et al., 2023a), instruction following (Jiang et al., 2023b), and value alignment (Liu et al., 2023a). Du et al. (2023) introduce a multi-agent debate framework where multiple LLM instances propose and debate their reasoning processes to reach a consensus, which improves the factuality and performance on arithmetic and strategic reasoning tasks. Similarly, ReConcile (Chen et al., 2023a) uses different backbone LLMs for debating and a confidence-weighted voting mechanism to obtain the final answer. Our approach differs in that we incorporate self-consistency estimates, and naturally calibrate the verbalized confidence through group deliberation with peer feedback and intermediate rationales for confidence adjustment, rather than manual rescaling in a post-hoc manner. We also incorporate both open and API-based models for the flexibility of supporting both logit-based and black-box confidence estimation. Liu et al. (2023b) propose an inference-time agent selection framework based on aggregated peer ratings over multiple rounds of interactions for a given query. In contrast, our agent selection process requires only the first-round individual confidence estimates at the task level.
@@ 3.1 Agent Ensemble and Stance Generation
To strive for a balance between task accuracy and calibration performance, it is necessary to maintain a certain level of diversity among the agents’ initial answers and reasoning paths while allocating the slots wisely so that the most suitable agents for the task can ideally become the majority, which serves as the basis for the ensemble. Inspired by Si et al. (2023b), we initialize four types of ”expert agents”. Each agent has a different prompting strategy or tool-use expertise: Chain-of-Thought (Wei et al., 2022) for multi-hop reasoning tasks, Program-of-Thoughts (Chen et al., 2022) for code and arithmetic reasoning tasks, Search-Augmented Self-Ask (Press et al., 2022) for factoid reasoning tasks, and GenRead prompting (Yu et al., 2023) for knowledge-intensive reasoning tasks. This initialization is flexible in that any new skill or prompting strategy can be easily added with modularity. As multiple skills might be relevant for an input dataset, we determine the importance ranking of each skill and accordingly allocate expert agents, based on a simple uncertainty-aware calibration score, detailed in Appendix A.1. We use Mistral-7B (Jiang et al., 2023a), GPT-3.5-turbo (OpenAI, 2022), and Cohere-Commend (Cohere, 2023) as backbones for the expert agents. For an incoming test question (e.g., ”Which element was discovered in 1898 and named after Greek ’new’?”), each expert agent performs self-deliberation by executing the corresponding prompting strategy and voting independently for an answer with a confidence estimate. The numerical confidence estimate can be based on the output sequence perplexity (PP) for open models  [^2] For open models with access to next-token probabilities, the raw confidence for the output sequence $W=(w_{1},...,w_{N})$ can be estimated as $\frac{1}{PP(W)}=P(w_{1},...,w_{N})^{1/N}$., or verbally elicited for black-box models, normalized to $[0,1]$. Note that this estimate is often unreliable and needs further calibration. We then obtain a set of unique and diverse answers (”stances”) by merging semantically equivalent pairs into clusters, using a GPT-3.5 judge following Tian et al. (2023). This constitutes the output of the first stage: semantically unique stances each with a corresponding frequency and aggregated mean confidence.
@@ A.4 Prompt templates and example output
We follow the prompt templates from the official implementations of CoT (Wei et al., 2022), PoT (Chen et al., 2022), and Knowledge prompting (Yu et al., 2023). For cost considerations, we adopt the zero-shot setting for each strategy (e.g. ”Let’s think/write Python programs step-by-step” without providing a demonstration) and resort to search verification only when the model is unsure about a generated premise. Other prompt templates, including the prompts for argument generation, feedback, and final confidence can be found in Table 2.